OpenStudy (anonymous):

help please!!!! Which equation would best help solve this problem? Five added to 3 times a number is equal to 8 less than five times the number. (Points : 1) 3x + 5 = 5x – 8 3x – 4 = 5x + 8 4x + 3 = 8x – 5 8x + 5 = 4x + 3

OpenStudy (anonymous):

A do you understand why?

OpenStudy (anonymous):

no im just lost on this one

OpenStudy (anonymous):

it is just a translation 5 is added to 3*x=5+3x 8 less thatn 5x=5x-8

OpenStudy (anonymous):

times is *

OpenStudy (anonymous):

@rapunzel206

OpenStudy (anonymous):

ohhhh i understand now i was a little confused but thank you for the help